OpenStudy (anonymous):

In basketball, the equation P=2x - x^2, describes a player's probability of scoring at least one point in a two shot free throw situation. P is the probability of scoring and x is the player's current free throw percentage. If Tony is a 40% free throw shooter what is the probability he will score at least one point the next time he is in a two shot situation?

OpenStudy (anonymous):

P = (2 * 0.4) - (0.4^2) = 0.64 There's a 64% chance that Tony will score at least one point the next time he is in a two shot situation.

OpenStudy (anonymous):

P=.8-16=15.2 how do you get 64%